If the pattern below were tripled, what would be the ratio of black squares to gray squares?


1
Expert's answer
2021-02-16T15:16:59-0500

number of black squares = 3

number of gray squares = 6

the ratio of the black squares to the gray squares is "=\frac{3}{6}=\frac{1}{2}"


on tripling the pattern,

number of black squares = 3*3 = 9

number of gray squares = 6*3 = 18

the ratio of the black squares and gray squares is = "\frac {9}{18}" = "\frac{1}{2}"

the ratio is same on tripling the pattern and equal to 1/2 .